OpenStudy (joannablackwelder):

Let f(x)=1/(x^2+1) and g(x)=x^-6. Find f(g(x)) and its domain.

OpenStudy (joannablackwelder):

I got 1/((x^-6)^2 + 1) which simplifies to x^3/(1+x^3)

OpenStudy (joannablackwelder):

That makes the domain all reals except for -1 the way I am looking at it, but that domain is not an option given.

OpenStudy (joannablackwelder):

@kropot72

OpenStudy (joannablackwelder):

The options for the domain are all real numbers x>0 x is not equal to 0 x is greater than or equal to 0

OpenStudy (joannablackwelder):

Is it all real numbers since the powers were originally even?

OpenStudy (joannablackwelder):

I thought the domain had to be real at each step though.

OpenStudy (freckles):

\[f(x)=\frac{1}{x^2+1} \text{ and we have } g(x)=x^{-6} \text{ aka } g(x)=\frac{1}{x^6}\]

OpenStudy (joannablackwelder):

Right.

OpenStudy (freckles):

so notice the domains of f and g separately first

OpenStudy (freckles):

the domain of f is all real numbers but what is the domain of g? -- and also I'm having trouble seeing how you got your f(g(x)) anyways

OpenStudy (joannablackwelder):

Domain of g is anything but 0

OpenStudy (freckles):

right @JoannaBlackwelder

OpenStudy (joannablackwelder):



OpenStudy (freckles):

so we already know we can't include 0 in the domain of f(g(x))

OpenStudy (freckles):

right so you should have